OpenStudy (anonymous):

y= sin 2x-x find y' (pi/2) a. -3 b. -1 c. 0 d. 1 e. undefined

OpenStudy (anonymous):

y=sin2x-x y'=2cos2x-1 y'=2cos(2*pie/2)-1 y'=2(-1)-1 y'=-3 (ans)

OpenStudy (anonymous):

why is there a 2 in front of cos2x

OpenStudy (anonymous):

coz it's a differentiation rule y=cosmx dy/dx=-msinmx